Mary has 6 times as much money as Bob. They are each given 6 dollars and now Mary has only 3 times as much money as Bob. How much money does each have before and after receiving the 6 dollars?
How can I solve this problem?

Answer by ankor@dixie-net.com(22740)

let m = Mary original money
let b = Bob's
:
Write an equation for each statement
:
"Mary has 6 times as much money as Bob."
m = 6b
:
They are each given 6 dollars and now Mary has only 3 times as much money as Bob.
m + 6 = 3(b+6)
m + 6 = 3b + 18
m = 3b + 18 - 6
m = 3b + 12
:
Now you can replace m with 6b, and find b, then find m
:


How much money does each have before and after receiving the 6 dollars?
How can I solve this problem?